Write slope-intercept form of the equation for the line passing through the given pair of points (-2,-6) and (9,2)
Answer by Alan3354(21600)

Write slope-intercept form of the equation for the line passing through the given pair of points (-2,-6) and (9,2)
-------------------------
1st find the slope, m
m = (y2-y1)/(x2-x1)
m = (2-(-6))/(9-(-2))
m = 8/11
Then use either point:
y-y1 = m*(x-x1)
y-2 = (8/11)*(x-9)
y-2 = 8x/11 - 72/11
y = (8/11)x - 50/11 slope-intercept form